OpenStudy (alexandervonhumboldt2):

What type of line is the graph of 4x + y = -3? (1 point) Vertical Horizontal Rising Falling D? due to bad english not sure about it

OpenStudy (anonymous):

Yes, it is falling.

OpenStudy (haleyelizabeth2017):

Correct :)

OpenStudy (anonymous):

https://www.desmos.com/calculator/4u754g7mzg

OpenStudy (haleyelizabeth2017):

All you have to do is subtract 4x from both sides and since it is a negative, it is a negative slope (falling) :)

jimthompson5910 (jim_thompson5910):

Solve for y 4x + y = -3 y = -3 - 4x y = -4x - 3 we see that the slope is -4. The negative slope means the line falls or goes downhill as you go from left to right.

OpenStudy (anonymous):

@AlexandervonHumboldt2 Since you have your answer, you can close your post. Thank you. :)